# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ Appendix C Error bounds for Matérn kernels
$\mathcal{M}_{\nu,\alpha}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)$
where $\Gamma$ denotes the gamma function and $K_{\nu}$ denotes the modified Bessel function of the second kind. Denoting the Euclidean distance of $x,y\in\mathbb{R}^{n}$ by $\|\cdot\|_{2}$, the Matérn kernel $k_{\mu,\alpha}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is given by
$k_{\nu,\alpha}(x,y)=\mathcal{M}_{\nu,\alpha}(\|x-y\|_{2}).$
The parameter $\nu$ controls the differentiability of the Matérn kernel at $x=y$ through $k_{\nu,\alpha}\in C^{2\lfloor\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ – with increasing $\nu$, the Matérn kernels become smoother. Equivalently, the sample paths of a Gaussian process with the Matérn kernel as covariance function are $\lfloor\nu\rfloor$-times mean-square differentiable, which provides a flexible way to control smoothness and makes Matérn kernels a popular choice in this setting. Another central property is the convergence
@@
to the squared exponential (Gaussian) kernel, uniformly on compact sets. Of particular interest are the cases $\nu=s+1/2$ for $s\in\mathbb{N}$: In that case, the Matérn covariance function is given as the product of a negative exponential and a polynomial of order $s$ [PorcBeviSchabOate2024]. Up to scaling constants, the Matérn kernel $k_{\nu,\alpha}$ on $\mathbb{R}^{n}$ induces the Sobolev space $H^{\nu+n/2}$ with equivalent norms; for the dimensionless Matérn kernel $k_{\nu}:=k_{\nu,\sqrt{2\nu}}$ the RKHS norm even coincides with the standard Sobolev norm of order $\nu+n/2$ up to a constant factor. We refer the reader to [FassYe11, PorcBeviSchabOate2024] for more insights into Matérn kernels.
Matérn kernels $k_{\nu}(x,y)=\mathcal{M}_{\nu,1}(\|x-y\|_{2})$ of smoothness $\nu>0$ induce the Sobolev spaces $H^{\nu+n/2}$. Denote the native norm of the RKHS by $\|f\|_{\mathbb{H}}$. The Matérn kernel is $2\lfloor\nu\rfloor$-times differentiable. Furthermore, (normalized) Matérn kernels are uniformly bounded by $1$. Thus, according to the discussion after [Wend04, Theorem 11.13], the bound
@@
Using the above properties, we obtain quasi-analogous results as in Section 4 for kernel EDMD approximations with underlying Matérn kernels. Following the derivations for Wendland kernels, we start with the deterministic error bound characterized by the projection error.
@@
Let $\nu>0$. Denote by $k_{\nu}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ the (dimensionless) Matérn kernel generating the Sobolev space $H^{\nu+n/2}(\Omega)$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\nu+n/2}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (23) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\nu+n/2$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (31) satisfies
$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}$ (56)
@@
The Matérn kernel satisfies $k_{\nu}\in C^{\lfloor 2\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. According to [Wend04, Theorem 11.13] and the subsequent discussion, we have
@@
Since the probabilistic error bound for the Matérn case is virtually identical to the Wendland case, merely inserting $\|k_{\nu}\|_{\infty}=1$, we state the analogue to Theorem 2.8 for the Matérn case below.
@@
Let the assumptions of Theorem 2.6 hold. Let $\nu>0$ and denote by $\mathbb{H}=H^{\nu+n/2}$ the reproducing kernel Hilbert space induced by the Matérn kernel $k_{\nu}$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}$, the full error admits the bound
@@
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$.
@@
the Matérn covariance function is monotonically decreasing for $r>0$, admitting its maximum at $\mathcal{M}_{\nu,\alpha}(0)=1$, so $\|k_{\nu}\|_{\infty}=1$. Similarly to Theorem 2.8, the result now follows immediately from Theorem 2.6. ∎
@@
for $\mathbb{H}=H^{\nu+n/2}$, following the line of argument in Proposition 4.5. To conclude, we state the analogue to the second part Corollary 4.12 for Matérn native spaces.
@@
Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for the RKHS $\mathbb{H}=H^{\nu+n/2}$ induced by the Matérn kernel $k_{\nu}$, we have the pointwise error bound
@@
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{\mathbb{H}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and constant $C>0$ only depending on $k_{\nu}$.
@@
The proof is analogous to the second part of Corollary 4.12 and leverages the uniform bound $\|k_{\nu}\|_{\infty}=1$. ∎
Appendix D Balancing between deterministic and probabilistic error
We discuss here the practical choice of the parameters $h_{\mathcal{X}}$, $m_{\mathrm{train}}$ and $\lambda$ for the stochastic kernel EDMD method when using the Wendland $C^{2}$ kernel ($l=1$) with spatial dimension $n=1$. The derivation for the Matérn kernel can be done in a similar fashion.
@@ Appendix D Balancing between deterministic and probabilistic error
from Corollary 4.3. Thus, decreasing fill distance improves the deterministic accuracy at rate $O(h_{\mathcal{X}}^{3/2})$. As $h_{\mathcal{X}}$ decreases, the number $d$ of design points increases, and the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ decreases. For the unregularized inverse, the bound $\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\leq\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}$ and $\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$ give
$\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X},\lambda}^{-1}v\right)^{1/2}=O\left(\sqrt{\frac{d}{d^{-3}}}\right)=O(d^{2})=O(h_{\mathcal{X}}^{-2}).$
@@
To prevent the regularization bias from dominating, one should require $\lambda=O(h_{\mathcal{X}}^{3})$.
Monte Carlo estimation of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ introduces a stochastic error. From Theorem 4.6, with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$,
$\|\widehat{\mathcal{K}}f-\widehat{\mathcal{K}}^{\mathrm{MC}}f\|_{\infty}\leq\varepsilon\,C_{\mathcal{X}}\|f\|_{\mathbb{H}}.$
Solving the probability expression for $\varepsilon$ gives
$\varepsilon=O(m^{-1/2}),$
again ignoring logarithmic terms in $d$. Since equispaced grids satisfy $d=O(h_{\mathcal{X}}^{-1})$, the Tikhonov control factor becomes
@@
$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3/2}\right),\quad\lambda(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3}\right)$

### v2

@@ Appendix C Error bounds for Matérn kernels
$\mathcal{M}_{\nu,\alpha}(r)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right)^{\nu}\text{\sf{k}}_{\nu}\left(\frac{\sqrt{2\nu}\,r}{\alpha}\right),$
where $\Gamma$ denotes the gamma function and $\text{\sf{k}}_{\nu}$ denotes the modified Bessel function of the second kind. Denoting the Euclidean distance of $x,y\in\mathbb{R}^{n}$ by $\|\cdot\|_{2}$, the Matérn kernel $\text{\sf{k}}_{\nu,\alpha}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is given by
$\text{\sf{k}}_{\nu,\alpha}(x,y)=\mathcal{M}_{\nu,\alpha}(\|x-y\|_{2}).$
The parameter $\nu$ controls the differentiability of the Matérn kernel at $x=y$ through $\text{\sf{k}}_{\nu,\alpha}\in C^{2\lfloor\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ – with increasing $\nu$, the Matérn kernels become smoother. Equivalently, the sample paths of a Gaussian process with the Matérn kernel as covariance function are $\lfloor\nu\rfloor$-times mean-square differentiable, which provides a flexible way to control smoothness and makes Matérn kernels a popular choice in this setting. Another central property is the convergence
@@
to the squared exponential (Gaussian) kernel, uniformly on compact sets. Of particular interest are the cases $\nu=s+1/2$ for $s\in\mathbb{N}$: In that case, the Matérn covariance function is given as the product of a negative exponential and a polynomial of order $s$ [42]. Up to scaling constants, the Matérn kernel $\text{\sf{k}}_{\nu,\alpha}$ on $\mathbb{R}^{n}$ induces the Sobolev space $H^{\nu+n/2}$ with equivalent norms; for the dimensionless Matérn kernel $\text{\sf{k}}_{\nu}:=\text{\sf{k}}_{\nu,\sqrt{2\nu}}$, the RKHS norm even coincides with the standard Sobolev norm of order $\nu+n/2$ up to a constant factor. We refer the reader to [14, 42] for more insights into Matérn kernels.
Matérn kernels $\text{\sf{k}}_{\nu}(x,y)=\mathcal{M}_{\nu,1}(\|x-y\|_{2})$ of smoothness $\nu>0$ induce the Sobolev spaces $H^{\nu+n/2}$. Denote the native norm of the RKHS by $\|\cdot\|_{\mathbb{H}}$. The Matérn kernel is $2\lfloor\nu\rfloor$-times differentiable. Furthermore, (normalized) Matérn kernels are uniformly bounded by $1$. Thus, according to the discussion after [50, Theorem 11.13], the bound
@@
Using the above properties, we obtain analogous results as in Section 4 for kernel EDMD approximations with underlying Matérn kernels. Following the derivations for Wendland kernels, we start with the deterministic error bound characterized by the projection error.
@@
Let $\nu>0$. Denote by $\text{\sf{k}}_{\nu}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ the (dimensionless) Matérn kernel generating the Sobolev space $H^{\nu+n/2}(\Omega)$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\nu+n/2}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (24) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\nu+n/2$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (32) satisfies
$\|\mathcal{K}-\widehat{\mathcal{K}}\|_{\mathbb{H}\to C_{b}}\leq C(1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}})\cdot h_{\mathcal{X}}^{\lfloor\nu\rfloor}.$ (57)
@@
The Matérn kernel satisfies $\text{\sf{k}}_{\nu}\in C^{\lfloor 2\nu\rfloor}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. According to [50, Theorem 11.13] and the subsequent discussion, we have
@@
Since the probabilistic error bound for the Matérn case is virtually identical to the Wendland case, merely inserting $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$, we state the analogue to Theorem 2.8 for the Matérn case below.
@@
Let the assumptions of Theorem 2.6 hold. Let $\nu>0$ and denote by $\mathbb{H}=H^{\nu+n/2}(\Omega)$ the reproducing kernel Hilbert space induced by the Matérn kernel $\text{\sf{k}}_{\nu}$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, the full error admits the bound
@@
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
@@
The Matérn covariance function is monotonically decreasing for $r>0$, admitting its maximum at $\mathcal{M}_{\nu,\alpha}(0)=1$, so $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$. Similarly to Theorem 2.8, the result now follows immediately from Theorem 2.6. ∎
@@
for $\mathbb{H}=H^{\nu+n/2}(\Omega)$, following the line of argument in Proposition 4.5. To conclude, we state the analogue to the second part Corollary 4.12 for Matérn native spaces.
@@
Let the assumptions of Theorem 2.6 hold and let $\lambda\geq 0$. Then, for the RKHS $\mathbb{H}=H^{\nu+n/2}(\Omega)$ induced by the Matérn kernel $\text{\sf{k}}_{\nu}$, we have the pointwise error bound
@@
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{\mathbb{H}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2}\right)$ and constant $C>0$ only depending on $\text{\sf{k}}_{\nu}$.
@@
The proof is analogous to the second part of Corollary 4.12 and leverages the uniform bound $\|\text{\sf{k}}_{\nu}\|_{\infty}=1$. ∎
Appendix D Balancing deterministic and probabilistic error
Here, we discuss the practical choice of the parameters $h_{\mathcal{X}}$, $m_{\mathrm{train}}$ and $\lambda$ for the stochastic kernel EDMD method when using the Wendland $C^{2}$-kernel ($l=1$) in spatial dimension $n=1$. The derivation for the Matérn kernel can be done in a similar fashion.
@@ Appendix D Balancing deterministic and probabilistic error
from Corollary 4.3. Thus, decreasing fill distance improves the deterministic accuracy at rate $O(h_{\mathcal{X}}^{3/2})$. As $h_{\mathcal{X}}$ decreases, the number $d$ of design points increases, and the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$ decreases. For the unregularized inverse, we start from the basic inequality $\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\leq\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}$. A classical result from kernel interpolation theory provides a lower bound on the smallest eigenvalue via the separation distance $q_{\mathcal{X}}=\min_{i\neq j}\|x_{i}-x_{j}\|_{2}$ of the set $\mathcal{X}=\{x_{1},\dots,x_{d}\}$: Specifically, we have $\lambda_{\min}(\mathbf{K}_{\mathcal{X}})\geq Cq_{\mathcal{X}}^{2s-n}$ with smoothness $s>n/2$ and dimension $n$ of the underlying kernel, see [51, 45, 50]. For Wendland kernels $\text{\sf{k}}_{n,l}$, where $s=\frac{n+1}{2}+l$, this gives rate $q_{\mathcal{X}}^{2l+1}$. We assume equispaced points in dimension $n=1$, for which $O(q_{\mathcal{X}})=O(h_{\mathcal{X}})$ holds, such that $\lambda_{\min}(\mathbf{K}_{\mathcal{X}})=O(q_{\mathcal{X}}^{3})=O(h_{\mathcal{X}}^{3})=O(d^{-3})$, and
$\left(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\right)^{1/2}=O\left(\sqrt{\frac{d}{d^{-3}}}\right)=O(d^{2})=O(h_{\mathcal{X}}^{-2}).$
@@
To prevent the regularization bias from dominating, one requires $\lambda=O(h_{\mathcal{X}}^{3})$.
Moreover, Monte Carlo estimation of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ introduces a stochastic error. From Theorem 4.6, with probability at least $1-2de^{-m\varepsilon^{2}/(2\|\text{\sf{k}}\|_{\infty})}$,
$\|\widehat{\mathcal{K}}f-\widehat{\mathcal{K}}^{\mathrm{MC}}f\|_{\infty}\leq\varepsilon\,C_{\mathcal{X}}\|f\|_{\mathbb{H}},$
where $C_{\mathcal{X}}$ depends on the data points $\mathcal{X}$. Solving the probability expression for $\varepsilon$ gives $\varepsilon=O(m^{-1/2})$, again ignoring logarithmic terms in $d$. Since equispaced grids satisfy $d=O(h_{\mathcal{X}}^{-1})$, the Tikhonov control factor becomes
@@
$m_{\mathrm{train}}(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3/2}\right),\quad\lambda(h_{\mathcal{X}})=O\!\left(h_{\mathcal{X}}^{-3}\right),$